OpenStudy (anonymous):

Please help! Find two functions, f and g, that will produce the given composite function, h=f o g. h(x)=(x+2)^2-5

OpenStudy (bahrom7893):

g(x)=x+2 f(x)=x^2-5

OpenStudy (bahrom7893):

f(g(x))=(x+2)^2-5

OpenStudy (anonymous):

Thanks bahrom! :)